OpenStudy (anonymous):

Evaluate. 3.2+1.4/7 --------- (9.3-7.6)4

OpenStudy (anonymous):

3.2 + 1.4/7 = 3.2 + 0.2 =3.4

OpenStudy (anonymous):

(9.3 - 7.6) x 4 = 1.7 x 4 = 6.8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :)